LBTO Technical/D/Service Night

OSA: S.Allanson

ISp: J.Power

Manager: M.Wagner

ISA: B. Rothberg

Eng. D.Ashby

=====================================

UT 01:30 - Enclosure opening

UT 01:52 - Moving to Rothberg target W2201+0226

UT 01:57 - Seeing ~1" average

UT 02:00 - preset failed due to winds

==========================================

UT 02:04 - Moving to V723 CAS (Local 1900 script)

UT 02:12 - would've been helpful if finder charts were made for each parallactic angle

UT 02:14 - starting spectra

UT 02:29 - airmass 1.131 seeing ~1"

UT 02:54 - MODS2 red got stuck on "1 of 3" but had taken two images - noted stuck on exposure 2, but it had completed exposure 2 (mods2r.20151125.0007.fits)

UT 03:01 - added an extra 10 minutes because of the third exposure not going forward.

UT 03:06 - 3rd red came out OK

UT 03:06 - moving to v339 Del

UT 03:17 - starting science

UT 03:22 - seeing 0.9" avg

UT 03:30 - Observations complete

============================================

UT 03:30 - trying W2201+0226 again

UT 03:44- starting science -conditions less than optimal - variable seeing and clouds rolling through

UT 04:08 - clouds have dispersed, mean seeing is 1.2" in guider 0.92" in WFS

UT 04:50 - finished target

====================================

UT 04:51 - moving to Feige110 - airmass 1.6

UT 04:58 - spectra started airmass 1.637

UT 05:09 completed, moving to G191B2b at airmass 1.28

UT 05:14 - tried to send preset for G191B2b, terminal hung (2x)

UT 05:15 - stopped and started MODS2 lbttcs - retried preset

UT 05:18 - noticed that after checking status the ISIS, Red and Blue Caliban on MODS2 data were off

UT 05:22 - started ISIS, red_caliban, blue_caliban, but xterm popped up with:

M2.CR% IN: IS>M2.CR ERROR: No Route to Destination Host M2.RC - host is unknown/unlisted
M2.CB% IN: IS>M2.CB ERROR: No Route to Destination Host M2.BC - host is unknown/unlisted

mods2data:mods% isisCmd m2.ie util heb_b status
*** ERROR: No Route to Destination Host M2.IE - host is unknown/unlisted

UT 05:25 - checked the KVM computers, both showed just the DOS prompt

UT 05:28 - still blank on the KVM - but MODS2 status shows everything running

=========================================================

UT 05:48 - switching to LUCI2

UT 06:10 - moving to M31 HK spectra - finder chart is terrible - will attempt acquisition

UT 06:19 - Too many stars, cannot find the actual target in the field image. I can recognize a few bright stars, but so many fainter stars, and without lines with distance measurements, hard to say.

=========================================================

UT 06:19 - Moving UDSK - LBTB

UT 06:40 - Unable to do MOS alignment - get error pop up when selecting LMS file:

Task: failed to draw LMS reference targets.

Message: No Matching LMS-targets coordinates found

Exception: NULL

=========================================================

UT 06:41 - moving to a longslit LUCI target - INAF DR Tau

UT 06:58 - science started - seeing 1.67"

UT 07:06 - DR Tau Imaging script

UT 07:13- DR Tau telluric

UT 07:26 - moving to VY Tau

UT 07:34 - science started

UT 07:37 - integration times too long (40sec DIT -fixing)

UT 07:42 - restarted script with new shorter DITS

UT 07:44 - LUCI2 stuck in grating and mask in FPU even though new script loaded - did an initialize all

UT 07:46 - Pointing is off checking now

UT 07:49 - skipped ahead and did not configure instrument

UT 07:52 - restarted scheduler

UT 08:04 - science started

UT 08:18 - Finished LUCI2 INAF spectra

===========================================

UT 08:19 - Moving to INAF Imaging Guaita H2imaging

UT 08:23 - no guidestar - checking pointing

UT 08:31 - corrected guidestar and made new script for H2 imaging

UT 09:22 - moving to Rubin-149 H2 std imaging

UT 09:25 - closing due to wall of clouds - guidestar for Rubin-149 (13-14th mag) not seen.

UT 09:51 - ran preset test with MODS2 - manual preset and acq script both worked

UT 09:57 - running pixel flats dual mode on MODS2 while we monitor all-sky - want to see if files save OK

UT 10:01 - data not showing up in /newdata or MODSDISP - ran fitsflush

UT 10:02 - data finally appeared in /newdata on mods1data computer and on MODSDISP

UT 10:03 - Opening up again

UT 10:09 - fluffier clouds rolling in from the south

UT 10:23 - trying Rubin-149/FS14 H2 again

UT 10:25 - Scheduler issue again, it skipped the instrument setup - restarting scheduler

UT 10:27 - resending star preset

UT 10:32 - 96 is last one don't use for this script

UT 10:34 - started a new script after abort and instrument and readout settings weren't updated

UT 10:35 - restarting scheduler

UT 10:36 - restarted script

UT 10:47 - going back to INAF Imaging Guaita H2imaging

UT 10:50 - guidestar not found checking pointing

UT 10:52 - restarting scheduler

UT 10:53 - sending preset

UT 11:16 - observer gui on obs2 doesn't show where we are in script - observer gui on luci machine does

UT 11:19 - clouds rolling in again, guide star fading in and out

UT 11:22 - lost the guidestar - aborting

UT 11:23 - switch to the LBCs - restarting scheduler as a precautionary

UT 11:25 - blinds filters in, n375 in, blind mask in

UT 11:35 - winds have picked up, clouds thick and threatening, closing up

UT 12:04 - dual pixelflats for MODS2

UT 12:24 - dual cal lamps for MODS2

UT 12:33 - slitflats - 0.8" and 1.2"

UT 13:09 - 8K Biases MODS2

-- %USERSIG{BarryRothberg - 2015-11-24}%

Comments

 
Topic revision: r4 - 25 Nov 2015, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback